The sum of three consecutive odd integers is 201. Find the integers.

Answer by checkley71(8403)

X (X=2) (X+4) ARE THE 3 CONSECUTIVE ODD NUMBERS.
X+(X+2)+(X+4)=201
3X+6=201
3X=201-6
3X=195
X=195/3
X=65 FOR THE SMALLEST ODD NUMBER
65+2=67 FOR THE MIDDLE ODD NUMBER.
65+4=69 FOR THE LARGEST ODD NUMBER.
PROOF
65+67+69=201
201=201